OpenStudy (anonymous):

Is -9, 1, 5,18 a answer to x-3=>9

OpenStudy (anonymous):

x = 3+9 =12

OpenStudy (anonymous):

if x-3 = 9

OpenStudy (anonymous):

THAT IS GREAT THEN OR EQUAL TO 9

OpenStudy (jit4won):

18 is the answer

OpenStudy (jit4won):

Just because x is greater than or equal to 12 and 18 is the best option from the given numbers